The weights of patients in a geriatric clinic are normally distributed with a mean of 60kg and a standard deviation of 10kg what is the probability that a patient picked at random has a weight between 50kg and 70kg


1
Expert's answer
2021-09-06T12:46:05-0400

Let "X=" weight of patient: "X\sim N(\mu, \sigma^2)."

Given "\mu=60\ kg, \sigma=10\ kg."


"P(50<X<70)=P(X<70)-P(X\leq50)"

"=P(Z<\dfrac{70-60}{10})-P(Z\leq\dfrac{50-60}{10})"

"=P(Z<1)-P(Z\leq-1)"

"\approx0.841344746-0.158655254"

"\approx0.6827"

By 68–95–99.7 rule


"P(50<X<70)=P(\mu-1\sigma<X<\mu+1\sigma)\approx"

"\approx0.6827"